Zelma is eighteen years older than her son. She was three times as old as he was one year ago. How old are they now?
Answer by Earlsdon(6294)

Let Z = Zelma's present age and S = her son's present age.
1) Z = S+18 "Zelma (Z) is 18 years older than her son (S)."
2) Z-1 = 3(S-1) "She was three times as old as he was one year ago."
How old are they now?
Rewrite equation 2) as:
2a) Z = 3(S-1)+1 Simplify and substitute into equation 1)
2b) Z = 3S-3+1
2c) Z = 3S-2
1a) 3S-2 = S+18
1b) 3S = S+20
1c) 2S = 20
1d) S = 10 This is her son's present age.
Z = S+18
Z = 10+18
Z = 28 This is Zelma's present age.